### v1

# TEMPO: Temporal Preference Optimization of Video LLMs via Difficulty Scheduling and Pre-SFT Alignment

Abstract
Video Large Language Models (Video LLMs) have achieved significant success by leveraging a two-stage paradigm: pretraining on large-scale video-text data for vision-language alignment, followed by supervised fine-tuning (SFT) for task-specific capabilities. However, existing approaches struggle with temporal reasoning due to weak temporal correspondence in the data and reliance on the next-token prediction paradigm during training. To address these limitations, we propose TEMPO (TEMporal Preference Optimization), a systematic framework that enhances Video LLMs’ temporal reasoning capabilities through Direct Preference Optimization (DPO). To facilitate this, we introduce an automated preference data generation pipeline that systematically constructs preference pairs by selecting videos that are rich in temporal information, designing video-specific perturbation strategies, and finally evaluating model responses on clean and perturbed video inputs. Our temporal alignment features two key innovations: curriculum learning which that progressively increases perturbation difficulty to improve model robustness and adaptability; and “Pre-SFT Alignment”, applying preference optimization before instruction tuning to prioritize fine-grained temporal comprehension. Extensive experiments demonstrate that our approach consistently improves Video LLM performance across multiple benchmarks with a relatively small set of self-generated DPO data. We further analyze the transferability of DPO data across architectures and the role of difficulty scheduling in optimization. Our findings highlight our TEMPO as a scalable and efficient complement to SFT-based methods, paving the way for developing reliable Video LLMs.
1 Introduction
Video LLMs have made significant strides in recent years, achieving remarkable success in various video understanding tasks [9, 22]. These advancements follow the standard multi-modal LLM paradigm: First, the model learns visual knowledge and vision-text alignment through large-scale pretraining on image/video-text pairs [19], and then it acquires instruction following capabilities via supervised fine-tuning (SFT) on curated instruction data [18, 29, 39]. This two-stage approach has enabled impressive improvements in video-language modeling, allowing models to generate coherent, contextually relevant responses conditioned on video input [8, 14, 26].
Despite this progress, current Video LLMs still exhibit poor temporal understanding, primarily due to two fundamental limitations. First, constructing high-quality video datasets is inherently challenging due to the complex visual and temporal structures in videos. Existing datasets often suffer from weak video-text correspondence, visual shortcuts, or insufficient temporal reasoning tasks [12, 21]. Second, both pretraining and SFT rely on next-token prediction using ground-truth text, which does not explicitly enforce dynamic temporal understanding. As a result, models frequently neglect subtle yet significant temporal details, over-rely on partial visual or textual cues, and struggle with temporal relationships between events, leading to inconsistent or flawed responses.
Figure 1: Example of detailed captioning result of the original Qwen2-VL-7B and our approach.
To address these limitations, in this paper, we propose TEMPO (TEMporal Preference Optimization), a systematic framework that tackles both challenges through: (1) a scalable, automatic dataset curation pipeline that generates high-quality temporal preference data, and (2) a tailored temporal alignment methodology that goes beyond next-token prediction. TEMPO is built on Direct Preference Optimization (DPO) [24], which optimizes the relative quality of model responses by maximizing the reward margin between preferred and non-preferred outputs. Unlike SFT, which requires high-quality, human-verified responses, DPO focuses on relative comparisons rather than absolute correctness—making it particularly effective for video tasks where perfect annotations are difficult to obtain. Subtle temporal aspects like sequence order and motion speed are critical for comprehensive video understanding yet challenging to annotate manually at scale. Our preference-based framework naturally discourages shortcut learning while enabling efficient scaling through automated annotation, making TEMPO an ideal methodology for enhancing temporal reasoning in Video LLMs.
We implement TEMPO through an automatic preference data generation pipeline that establishes the foundation for effective temporal alignment in Video LLMs. Inspired by advances in image-based LLMs [33, 41], we generate preference pairs by evaluating model responses under both clean and corrupted visual inputs. Our pipeline consists of three key stages: (1) curating a diverse set of videos that emphasize complex temporal relationships; (2) introducing targeted temporal perturbations—including frame dropping, sequence reversal, and frame shuffling—that precisely target known temporal reasoning weaknesses of current Video LLMs; and (3) constructing preference pairs by contrasting the model’s outputs for clean versus temporally corrupted videos. We also introduce a difficulty factor that quantifies perturbation magnitude for controlling the sample difficulties. This systematic approach enables us to efficiently generate high-quality contrastive examples that directly highlight temporal understanding capabilities without requiring expensive human annotation.
Our tailored temporal alignment framework consists of two key technical innovations. First, we implement curriculum learning with progressive difficulty lifting, which gradually increases sample complexity during optimization. This controlled exposure to increasingly challenging temporal constrastive pairs enables more effective learning. Second, we introduce a novel “Pre-SFT Alignment” strategy, which reverses the conventional order of model alignment. Traditional approaches apply DPO after instruction-tuning (SFT) [30], assuming the model has already learned basic capabilities. In contrast, we argue that robust temporal understanding should be established first, before learning to follow diverse instructions. Our Pre-SFT Alignment applies DPO first on our curated dataset, forcing the model to develop strong temporal reasoning and video-text correspondence prior to any instruction tuning.
Through extensive evaluation, we demonstrate that TEMPO consistently improves model performance across multiple benchmarks including Video-MME [8], MLVU [40] and Vinoground [36] with only a small number of self-generated DPO pairs—eliminating dependence on external models or APIs. Our ablation studies validate the effectiveness of both technical contributions: the difficulty curriculum significantly stabilizes training and improves final performance, while the Pre-SFT Alignment strategy significantly outperforms the conventional SFT-then-DPO approach, particularly on tasks requiring fine-grained temporal reasoning. Further analysis reveals valuable insights about DPO data transferability between different model architectures and scales, demonstrating the framework’s flexibility. By introducing this automated, structured preference learning pipeline, we advance the state of video LLMs and pave the way for more robust and temporally-aware multi-modal models.
Figure 2: Illustration of our DPO data generation pipeline.
2 Method
In this section, we first present a preliminary analysis on the detailed captioning behavior of current Video LLMs, highlighting the need for a more effective alignment strategy. We then introduce our DPO-based approach and propose a three-stage pipeline for automatically generating preference data by comparing model responses on clean and perturbed video inputs. As part of this process, we incorporate a difficulty factor to control the magnitude of noise, enabling structured learning. Finally, we present our novel “Pre-SFT Alignment” paradigm, where we apply DPO using captioning-based preference pairs before SFT, enhancing fine-grained temporal alignment in Video LLMs.
2.1 Preliminary Analysis on Detailed Captioning
We conduct a preliminary analysis of the detailed captioning behavior of Qwen2-VL-7B [31] by manually annotating events in video keyframes and comparing them to events captured in model-generated captions across 19 videos. The comparison reveals a precision of 74.4% and a recall of merely 35.0%, indicating that the model not only introduces a considerable amount of hallucinated content but also overlooks a substantial portion of events present in the videos (example shown in Figure 1). These findings underscore the critical need for improving video LLMs’ ability to capture and retain temporal information accurately.
2.2 DPO Data Generation Pipeline
Video DPO data consists of four key components: video, instruction, chosen response, and rejected response. Inspired by POVID [41], we generate video DPO data by feeding clean and noised visual inputs into Video LLMs, using the response to the clean input as the chosen response and the response to the corrupted input as the rejected response. To enable preference learning without reliance on external LLMs or APIs, we design a structured pipeline for generating video DPO data, which follows a three-stage process to systematically construct videos, craft instructions, and derive response pairs (Figure 2).
2.2.1 Video Selection
Public video datasets often contain relatively static videos with minimal variation in visual content, making them easy for models to process but ineffective for learning temporal structure. Such videos provide limited benefit for DPO and are excluded from our selection process. Instead, we prioritize videos featuring multiple, well-separated events with distinct yet connected visual content. To achieve this, we implement a three-step filtering procedure to identify suitable videos and extract clear keyframes.
In the first step, we perform scene boundary detection using TransNetV2 [27], which offers more accurate segmentation than tools like PySceneDetect [^1] https://github.com/Breakthrough/PySceneDetect, particularly in handling gradual transitions such as fade-ins and fade-outs. The number of detected scenes serves as an indicator of a video’s temporal complexity, helping us distinguish between dynamic and static content. To refine our selection, we discard monochrome scenes shorter than 0.2 seconds, as these typically correspond to transition frames with little meaningful information. We also filter out videos containing scenes longer than 16 seconds, as such segments often consist of unstructured long shots that are difficult to decompose into discrete events, making them unsuitable for preference learning. Since current segmentation tools struggle to effectively break down these complex videos, we discard them in favor of those with clearer event boundaries.
In the second step, we further refine our selection by filtering out videos that contain frequent scene cuts but little new visual information. This includes cases where multiple shots capture the same scene from different angles or repetitive back-and-forth cuts, such as those commonly found in interview footage. To achieve this, we introduce a similarity grouping stage, which computes the cosine similarity between middle frames of clips using SigLIP [35] and iteratively group clips with similarity above a predefined threshold. If a video contains too few unique groups, it is deemed to have low information density and is removed. Conversely, if a video contains too many unique groups, it is considered excessively complex, presenting challenges for both computational efficiency and model capability. To balance these factors, we retain only videos with 4 to 32 distinct groups, ensuring sufficient diversity while maintaining manageable complexity.
In the final step, we extract keyframes that best represent the visual content of each selected clip. Since we have limited clip length to a maximum of 16 seconds, ensuring that each clip corresponds to a single event, we choose to extract two keyframes per clip, which is generally sufficient to capture its core elements. To improve the clarity of the selected keyframes, we set anchor points at one-third and two-thirds of the clip’s duration and select frames that are both close to these anchors and exhibit high visual clarity. Clarity is assessed using the Laplacian operator [11], which measures sharpness to filter out blurry frames. This final step ensures that the extracted keyframes provide distinct and representative snapshots of the video, further enhancing the quality of our DPO data.
Through this three-step process, we systematically filter videos to maximize temporal diversity, eliminate redundant or overly simplistic content, and ensure that selected clips contain clear, well-defined events. This careful curation allows us to construct a dataset that effectively supports Video LLMs in learning robust temporal understanding through preference optimization.
2.2.2 Perturbation Strategy and Task Design
The design of perturbation strategies and task formulation is critical in our DPO data construction pipeline, as they directly influence the model behaviors we aim to reinforce. Given that video-based QA datasets often suffer from low quality and unreliable annotations, we opt to use detailed captioning as the sole task for preference learning. By focusing on captioning rather than QA, we ensure that preference signals are derived from a more structured and reliable task formulation.
To introduce meaningful preference distinctions, we carefully design perturbation strategies that target the known weaknesses of Video LLMs. One major issue is the model’s tendency to overlook important details in videos or rely on superficial shortcuts in visual or textual inputs. To counter this, we introduce random clip dropping, where certain clips are removed from the video before being fed into the model. This forces the model to generate a response with missing information, leading to a rejected response that either omits critical details or makes unreliable inferences based on partial context. This perturbation explicitly penalizes shortcut reliance and encourages the model to develop a more comprehensive understanding of the input.
Another common failure mode in Video LLMs is the difficulty in maintaining temporal structure and distinguishing between visually similar but sequentially distinct frames. To address this, we introduce random clip shuffling and clip reversal, where clips within a video are either randomly reordered or flipped in temporal sequence. Unlike clip dropping, this perturbation preserves the total amount of visual information but disrupts its logical arrangement. Consequently, the model produces a rejected response that misinterprets event chronology or blends together unrelated segments, highlighting its struggle with temporal reasoning. By optimizing for preference under this perturbation, we encourage the model to develop stronger temporal comprehension and robustness against structural inconsistencies.
These perturbation strategies serve as structured adversarial challenges, enabling the model to learn more precise and reliable video understanding through preference optimization. By constructing preference pairs that reflect real-world failure cases, we ensure that DPO training directly addresses the fundamental shortcomings of current Video LLMs.
2.2.3 Response Generation
Generating high-quality, detailed captions for videos with multiple events is a challenging task for Video LLMs, as they often struggle to retain and integrate information across extended sequences. To mitigate this issue, we draw inspiration from prior work in video captioning and adopt a contextual captioning strategy that enables the model to capture fine-grained temporal dependencies without overwhelming its processing capacity.
Instead of requesting a single caption for the entire video, we employ a contextualized clip captioning approach [3]. Specifically, at each step, we provide the model with two consecutive clips (corresponding to four frames) and prompt it to generate a caption for the latter clip while considering the contextual information from both. This localized approach ensures that captions remain detailed while still maintaining continuity across the video. This strategy effectively injects contextual dependencies into each caption while preventing information overload.
Once individual clip captions are generated, we aggregate them to form a global video-level caption. To achieve this, we feed all clip captions into the model and prompt it to generate a comprehensive, structured summary of the entire video. This multi-stage process ensures that the final caption preserves both fine-grained event details and high-level narrative coherence.
For response generation in the clean input setting, we directly use the model’s output given the full sequence of clip captions. For perturbed inputs, we modify the clip captions according to our designed perturbations—for instance, dropping, shuffling, or reversing some of the clips. The model is then prompted with the altered captions to generate a corresponding response, which serves as the rejected response in our preference learning pipeline. By structuring response generation in this manner, we ensure that our preference pairs reflect meaningful variations in video comprehension, reinforcing the model’s ability to process complex temporal relationships and mitigate common failure cases.
2.3 Difficulty Scheduling
Figure 3: Illustration of video perturbation strategies controlled by the difficulty factor $r$.
To systematically examine the impact of perturbation difficulty in DPO, we introduce a difficulty factor $r$ that controls the magnitude of noise applied during preference data generation. This parameter enables us to adjust the degree of disruption in a fine-grained manner, ensuring a structured and progressive learning process for the model.
Given a video with $N$ clips, we regulate difficulty as follows: (i) For clip dropping, we retain only $\lceil N/r\rceil$ clips while discarding the rest, reducing the available visual information. (ii) For clip shuffling and reversal, we partition the clips into $\lceil N/r\rceil$ groups, treating each group as an indivisible unit during perturbation. This ensures that lower values of $r$ introduce finer-grained perturbations while preserving more of the original structure, whereas higher values of $r$ cause more drastic disruptions.
Figure 3 shows an illustration of how the videos are perturbed for different difficulty factors. Intuitively, a smaller $r$ results in weaker perturbations, meaning more information and temporal consistency are preserved, making it harder to differentiate between clean and perturbed inputs. Conversely, a larger $r$ leads to stronger perturbations, making distinctions more apparent and simplifying the preference learning task.
Based on this difficulty factor, we adopt a curriculum learning strategy, where $r$ is gradually decreased throughout training. This allows the model to first learn from easier cases with clearer distinctions and progressively refine its ability to handle more subtle perturbations. By increasing difficulty over time, we ensure that the model develops a more nuanced understanding of video content, leading to improved alignment with fine-grained temporal structures.
2.4 Pre-SFT Alignment
Figure 4: Training loss and gradient norm during SFT. Pre-SFT alignment leads to lower loss and stable gradients during training.
In the conventional training paradigm, DPO is typically applied after SFT. This ordering ensures that the model first acquires the fundamental ability to perform tasks and follow instructions before refining its alignment with human preferences. However, considering our objective of improving the model’s fine-grained understanding of video content and its temporal structure, we propose an alternative strategy: performing DPO before SFT. Our key insight is that the primary limitation of Video LLMs lies in their inability to establish precise correspondences between visual input and textual descriptions. Without a strong foundation in fine-grained video-text alignment, the model is prone to relying on shortcuts and generating responses that are only loosely connected to the visual content. This issue is particularly pronounced in video-based tasks, where temporal dependencies play a crucial role in determining meaning. By applying DPO before SFT, we enable the model to first internalize the intricate relationships between video frames and textual descriptions before moving on to more general instruction-following training. As shown in Figure 4, we observe that models trained with Pre-SFT Alignment exhibit significantly lower loss values and more stable gradient norms during the subsequent SFT phase. This suggests that early-stage temporal alignment enhances the model’s ability to process visual information effectively, leading to improved convergence and overall training stability. By shifting alignment to an earlier stage in training, our approach ensures that Video LLMs develop a robust foundation in temporal understanding, making them better equipped to handle complex multi-modal reasoning tasks.
3 Experiment
#Params VideoMME MLVU Vinoground
TP TR Overall Overall Text Video Group Average
Closed-source Models
Gemini-1.5-pro [9] - - - 75.0 - 35.8 22.6 10.2 22.9
GPT-4o [22] - - - 77.2 64.6 54.0 38.2 24.6 38.9
Claude 3.5 Sonnet [1] - - - 60.0 - 32.8 28.8 10.6 24.1
Open-source Models
Qwen2-VL [31] 72B - - 71.2 - 50.4 32.6 17.4 33.5
VideoLLaMA2 [6] 72B - - 62.4 61.2 36.2 21.8 8.4 22.1
LLaVA-Video [39] 7B - - 63.3 70.8 - - - -
LLaVA-OneVision [13] 7B - - 58.2 64.7 41.6 29.4 14.6 28.5
MiniCPM-V-2.6 [5] 8B - - 60.9 - 32.6 29.2 11.2 24.3
Video-LLaVA [16] 7B - - 39.9 47.3 24.8 25.8 6.6 19.1
Qwen2-VL-2B 2B 54.5 34.5 55.2 57.7 26.4 24.6 6.2 19.1
   + SFT 2B 65.5 32.8 54.6 60.0 27.4 26.0 5.4 19.6
   + TEMPO (Ours) 2B 63.6  (+9.1) 35.0  (+0.5) 55.1 (-0.1) 60.2  (+2.5) 28.0  (+1.6) 25.2  (+0.6) 5.2 (-1.0) 19.5  (+0.4)
Qwen2-VL-7B 7B 70.9 42.4 60.4 64.5 34.0 26.6 10.6 23.7
   + SFT 7B 70.9 44.1 60.7 64.6 34.6 26.6 9.6 23.6
   + TEMPO (Ours) 7B 74.5  (+3.6) 45.2  (+2.8) 61.0  (+0.6) 65.8  (+1.3) 37.2  (+3.2) 29.6  (+3.0) 12.2  (+1.6) 26.3  (+2.6)
Table 1: Benchmark results of our approach on VideoMME, MLVU, and Vinoground. TP and TR denotes Temporal Perception and Temporal Reasoning, respectively, two subtasks of VideoMME.
3.1 Setting
Original After Step 1 After Step 2
YouTube 1-2m 22427 8724 159
YouTube 2-3m 24685 2191 597
Total 47112 10915 2195
Table 2: The number of remaining videos after each step of our data pipeline.
We conduct our main experiments using the Qwen2-VL model family [31], focusing on two model scales: 2B and 7B parameters, due to its strong performance in open-source video LLMs. During data construction, the videos are sourced from a subset of the LLaVA-Video-178K dataset [39], i.e., Youtube videos within the duration of 1-3 minutes. After each step of our data pipeline, the number of remaining videos is shown in Table 2. To systematically analyze the impact of perturbation difficulty, we introduce four levels of difficulty, with the value of $r$ as 2, 4, 8, and 16. The final data contains 6,385 training samples for each difficulty level, resulting in a total of 25,540 examples.
For DPO training, we adopt a four-stage progressive difficulty schedule, where the model is trained sequentially on increasingly challenging data. Beginning with the easiest perturbation setting ($r=16$), the model gradually adapts to more complex data at $r=8$, $r=4$, and finally $r=2$. At each stage, only the data corresponding to that specific difficulty is used. To ensure efficient training, we apply LoRA [10] with rank 32, $\alpha=64$, a batch size of 16, and a learning rate of 5e-6, while keeping the ViT module frozen. Each training stage lasts for approximately one hour on 8 NVIDIA H800 GPUs, accumulating to a total of four hours for the entire DPO process.
For SFT, we use a dataset of 60,000 samples randomly drawn from LLaVA-Video-178K [39]. Unlike in DPO, the ViT module is unfrozen to allow for additional adaptation. The SFT phase follows a similar training setup with LoRA rank set to 32, $\alpha$ set to 32, a batch size of 16, and a higher learning rate of 1e-4. The entire fine-tuning process takes approximately three hours on the same hardware setup.
Both DPO and SFT follow the original Qwen2-VL preprocessing pipeline, where videos are sampled at 2 FPS. While extracted keyframes are used for data construction, they are not used during training. To accommodate memory constraints, we limit the number of pixels per frame to 90,000, the maximum number of frames per video to 100, and the total sequence length to 8,192.
Evaluation Datasets
For evaluation, we assess models trained with and without DPO on three widely used benchmarks. VideoMME [8] serves as a comprehensive benchmark for general video-language understanding, MLVU [40] focuses on long-video comprehension, and Vinoground [36] specifically evaluates fine-grained temporal reasoning. To ensure high-quality video processing during inference, we increase the maximum pixels per frame to 250,000 and limit the number of frames per video to 64.
3.2 Main Results
We evaluate the effectiveness of our method by comparing the performance of three models: the original model without additional tuning, the model trained directly with supervised fine-tuning (SFT), and the model first tuned on our DPO data before undergoing SFT (Ours). The results, presented in Table 1, highlight the impact of our approach across multiple benchmarks.
Our method consistently outperforms both the original and SFT-trained models, demonstrating its ability to enhance temporal understanding in video-language models. Notably, on VideoMME, our model achieves substantial improvements on the temporal-related subtasks, with a gain of +3.6 points in Temporal Perception and +2.8 points in Temporal Reasoning. This suggests that the fine-grained alignment introduced by our DPO framework effectively reinforces the model’s capacity to process and reason about temporal information. Furthermore, our method yields a significant boost on Vinoground, a benchmark specifically designed to assess fine-grained temporal understanding, improving performance by 2.6 points—representing a relative increase of over 10%.
These results provide strong empirical evidence that our approach successfully enhances video LLMs by improving their ability to capture temporal structure. By incorporating fine-grained alignment before SFT, our method injects a more robust temporal understanding into the model, leading to consistent gains across diverse video-language tasks.
3.3 Analysis
Transferrability across models
VideoMME MLVU Vinoground
Overall Overall Text Video Group Avg
Qwen2.5-VL-7B 63.1 64.0 41.6 34.6 17.6 31.3
   + SFT 63.1 65.5 42.8 28.8 14.6 28.7
   + TEMPO 63.8 67.3 44.2 25.8 12.4 27.5
InternVL2.5-8B 64.4 67.8 42.6 23.8 10.2 25.5
   + SFT 64.7 67.4 42.8 24.4 10.2 25.8
   + TEMPO 64.3 67.9 45.0 24.4 10.4 26.6
Table 3: Experiment results on the transferrability of our approach across model architectures.
Figure 5: Performance comparison between different tuning strategies. Our TEMPO brings significant boost for downstream tasks by enhancing the temporal alignment before SFT.
In our main experiment, we adopt a self-generation approach, using the same model for both generating DPO data and performing DPO. This design choice is crucial, as it ensures that the perturbations and response preferences align closely with the model’s inherent strengths and weaknesses. To assess the transferability of our DPO data across different architectures, we conduct an additional experiment where the DPO data generated by Qwen2-VL-7B is used to train two other models: Qwen2.5-VL-7B [2], which shares a similar architecture, and InternVL2.5-8B [4], which differs more significantly. The results, presented in Table 3, show that applying DPO on these models using externally generated data still yields improvements over standard SFT on certain metrics, indicating some level of transferability. However, the gains are neither as consistent nor as pronounced as those observed with self-generated data. This suggests that while DPO data can generalize to some extent, aligning the data generation process with the target model leads to more effective learning, likely due to a better match in learned representations and response tendencies.
VideoMME MLVU Vinoground
Overall Overall Text Video Group Avg
Qwen2-VL-7B 60.4 64.5 34 26.6 10.6 23.7
   + SFT 60.7 64.6 34.6 26.6 9.6 23.6
   + TEMPO (2B-version data) 60.9 64.3 37.4 28.2 12.4 26.0
Qwen2-VL-2B 55.2 57.7 26.4 24.6 6.2 19.1
   + SFT 54.6 60.0 27.4 26.0 5.4 19.6
   + TEMPO (7B-version data) 52.9 59.5 26.0 24.6 6.2 18.9
Table 4: Experiment results on the transferrability of our approach across model sizes.
Figure 6: Vinoground score for models trained with DPO data of different difficulty levels.
Transferrability across model sizes
We further investigate the transferability of our approach across model sizes. Specifically, we evaluate the impact of tuning a larger model using DPO data generated by a smaller model and vice versa. As shown in Table 4, when applying the DPO data from Qwen2-VL-2B to fine-tune Qwen2-VL-7B, the model demonstrates improvements on Vinoground, slight gains on VideoMME, but a minor performance drop on MLVU compared to the SFT-only baseline. In contrast, when tuning Qwen2-VL-2B with DPO data generated by Qwen2-VL-7B, the model consistently underperforms relative to the SFT model. These results emphasize the importance of self-generated DPO data, as it better aligns with the model’s specific capabilities and learning patterns. The findings suggest that while some benefits can be retained when transferring data across sizes, DPO is most effective when the data is tailored to the model’s own behavior and representations.
Ablation Study on Pre-SFT Alignment
To showcase the effectiveness of the Pre-SFT Alignment strategy, we conduct an ablation study comparing the performance of multiple different tuning orders. Figure 5 highlights the performance differences of Qwen2-VL-7B, comparing its results after tuning with various strategies against the original base model. The Pre-SFT alignment strategy consistently outperforms all other approaches across all three benchmarks. This demonstrates its ability to effectively combine the benefits of both realms: enhancing the model’s temporal understanding while preserving high levels of instruction-following capability. This ablation study underscores the advantages of aligning the model with fine-grained temporal knowledge before instruction fine-tuning, enabling better overall performance.
Analysis of Difficulty Scheduling
To investigate the impact of DPO data difficulty on model performance, we first conduct a preliminary comparison by tuning Qwen2-VL-7B with a single split of DPO data containing the same difficulty level and evaluating the models on Vinoground. As shown by the bars in Figure 6, models tuned with easier DPO data exhibit better performance. This suggests that, due to the initial inability of the model to capture temporal structure, preference optimization with easier data proves more beneficial in the early stages of alignment. This observation motivates our curriculum learning strategy, which formulates DPO as a staged process, progressively increasing the difficulty of DPO data throughout the course. As shown by the line in Figure 6, this progressive tuning approach leads to steady improvements in model performance, validating the efficacy of our strategy in helping the model gradually refine its temporal understanding.
4 Related Work
Video Large Languge Models
Building upon early efforts to integrate visual understanding into powerful language models [19, 18, 42], current Video LLMs typically consist of three key components: a visual encoder, a multimodal projector, and a language model [28]. The visual encoder processes video frames into patchified feature sequences, which are then transformed by the multimodal projector into discrete visual tokens. These visual tokens, along with text tokens, serve as inputs to the language model, enabling it to generate textual responses conditioned on video content. Efforts to improve video LLMs have largely followed two main directions. One approach focuses on expanding the scale and improving the quality of video datasets [39, 3, 4], often through the use of external annotation tools and proprietary LLMs to generate detailed captions. Another direction aims to extend the applicability of Video LLMs by enabling them to handle long-duration videos [37, 26, 25], high-resolution content [31, 2, 5, 13], or streaming video inputs [17]. In contrast, our work addresses a more fundamental challenge in video LLMs: the acquisition of temporal knowledge. Prior studies have frequently highlighted the deficiency of these models in capturing fine-grained temporal structure, which we directly tackle through our proposed approach.
Direct Preference Optimization
Originally introduced in the text domain as an alternative to Reinforcement Learning from Human Feedback (RLHF) [23], Direct Preference Optimization (DPO) [24] was first employed to align language model behavior with human preferences, particularly to mitigate the generation of harmful or undesirable responses. DPO optimizes an implicit reward model derived from preference data, providing a straightforward and effective approach to post-training alignment. Due to its simplicity and strong performance, DPO has gained widespread adoption, leading to various improvements such as KTO [7] and $\beta$-DPO [32]. More recently, research has explored the adaptation of DPO and other alignment strategies to multi-modal LLMs by developing carefully curated data generation pipelines. For instance, POVID [41] constructs preference data for image-language models by comparing responses to clean and noised visual inputs, while MIA-DPO [20] generates preference data for multi-image understanding by concatenating single-image datasets and assessing whether model responses correctly attend to the intended image. Along with some concurrent work [38, 15, 34], we take a pioneering step toward leveraging DPO for improving the temporal reasoning capabilities of Video LLMs, addressing a key limitation that has been frequently highlighted in prior studies.
5 Conclusion
In this work, we propose a novel approach to enhance the temporal understanding of video LLMs through DPO, addressing the deficiency in capturing fine-grained temporal structures in current models. Our method leverages a carefully designed data generation pipeline to curate video preference data and a novel curriculum learning + Pre-SFT Alignment strategy to inject temporal understanding abilities into video LLMs. We demonstrate the effectiveness of our approach through extensive experiments on multiple benchmarks, showing substantial gains in temporal reasoning and perception. These results open the door for further improvements in video LLMs, particularly in refining their ability to understand and reason about temporal dynamics in videos.